# Estimation of Out-of-Sample Sharpe Ratio for High Dimensional Portfolio Optimization

### v1

@@ 4 Numerical Experiments
In this section, we perform numerical experiments to validate our theoretical findings. In Section 4.1, we provide the basic settings for SR estimation and verify Theorem 2.7, especially the asymptotic behavior as $n$ increases. In Section 4.2, we verify our estimation of the efficient frontier in Theorem 3.2 as we change target portfolio returns. Some additional simulation settings are presented in Appendix A to provide more comparisons on different choices of the population covariance ${\bm{\Sigma}}$, the mean vector of returns ${\bm{\mu}}$ and the regularization term $\mathbf{Q}$. In all the simulations, we set $r_{0}=0$ for simplicity and in this case, $\mathbf{r}={\bm{\mu}}$.
@@ 4.1 Sharpe ratio estimation
We present the simulation results under the basic settings. We first fix the sample size as $n=1500$ and consider two dimensions: $p=750$ (ratio $c=1/2$) and $p=2250$ (ratio $c=3/2$). These cases correspond to values of $c$ less than and greater than 1. To simulate financial returns, we construct the population covariance matrix ${\bm{\Sigma}}_{0}$, mean vector ${\bm{\mu}}_{0}$, and the matrix $\mathbf{Q}_{0}$ as follows:
@@
    The population covariance ${\bm{\Sigma}}={\bm{\Sigma}}_{0}={\rm diag}(\lambda_{1},\dots,\lambda_{p})+2{\bm{1}}{\bm{1}}^{\top}$, where $\{\lambda_{i}\}_{i=1}^{p}$ are generated from a truncated $\Gamma^{-1}(1,1)$ distribution, truncated with the interval $[0.01,9]$, and then ranked in decreasing order. This particular construction of ${\bm{\Sigma}}$ follows from the factor model, mimicking the well-accepted low-rank plus sparse structure of the covariance matrix of financial returns. Specifically, the covariance consists of the market systematic risk, where the vector ${\bm{1}}$ represents a market factor introducing a spike in covariance spectrum, and the specific risk, where $\lambda_{i}$ represents the individual risk of each asset.
@@
    The regularization matrix is $\mathbf{Q}=q\cdot\mathbf{Q}_{0}$. The matrix $\mathbf{Q}_{0}$ is defined as ${\rm diag}(3,...,3,1,...,1)$, where the numbers of $3$ and $1$ entries are both $p/2$. By structuring $\mathbf{Q}_{0}$ in this manner, we categorize stocks into two distinct volatility regimes, high and low, roughly capturing some weak knowledge of the covariance structure. In the simulation, we will vary the values of $q$ in order to better verify Theorem 2.7.
In the simulation, for each pair of $(n,p)$ considered, the parameters ${\bm{\Sigma}}_{0}\in\mathbb{R}^{p\times p}$, ${\bm{\mu}}_{0}\in\mathbb{R}^{p}$, and $\mathbf{Q}_{0}\in\mathbb{R}^{p\times p}$ are generated. Once they are generated, they remain fixed throughout the entire simulation procedure. Then we generate i.i.d. Gaussian random vector $\mathbf{R}_{i}$ ($i\in[n]$) with mean ${\bm{\mu}}={\bm{\mu}}_{0}$ and variance ${\bm{\Sigma}}={\bm{\Sigma}}_{0}$. The observed random matrix is $\mathbf{R}=(\mathbf{R}_{1},\mathbf{R}_{2},...,\mathbf{R}_{n})^{\top}\in\mathbb{R}^{n\times p}$. Here, we consider $(n,p)=(1500,750)$ and $(n,p)=(1500,2250)$. Note that since we assume the mean vector ${\bm{\mu}}$ is known, $\mathbf{X}=\mathbf{R}-{\bm{1}}_{n}{\bm{\mu}}_{0}^{\top}$ is also well observed. Recall that $\widehat{\bm{\Sigma}}=\mathbf{X}^{\top}\mathbf{X}/n$, and
We compare the true $SR(\mathbf{Q})$ and the estimator $\widehat{SR}(\mathbf{Q})$ in Theorem 2.7 which is
@@
The Sharpe computation is then carried out as follows for each pair of $(n,p)$. We repeat the generation of Gaussian random matrices $\mathbf{R}\in\mathbb{R}^{n\times p}$ for $100$ independent times. For each repetition, we vary the value of $q$ in $\mathbf{Q}$ to compute $SR(q\cdot\mathbf{Q}_{0})$ and $\widehat{SR}(q\cdot\mathbf{Q}_{0})$. If $c<1$, we vary $q$ from (1:30)/5(1:30)/5( 1 : 30 ) / 5, while if $c>1$, we vary $q$ from (1:30)/1.5(1:30)/1.5( 1 : 30 ) / 1.5. Here, we assess the alignment between $SR(\mathbf{Q})$ and the proposed statistics $\widehat{SR}(\mathbf{Q})$ with various $q$ values. For each given $q$, we will obtain 100 different values of $SR(q\cdot\mathbf{Q}_{0})$ and $\widehat{SR}(q\cdot\mathbf{Q}_{0})$.
In Figure 1, the $x$-axis represents the different values of $q$, while the $y$-axis gives the Sharpe ratio. The simulation results are presented as the average of $SR(q\cdot\mathbf{Q}_{0})$ and $\widehat{SR}(q\cdot\mathbf{Q}_{0})$ over $100$ independent trials.
Figure 1: Simulation results in the basic settings. Figure 1(a) shows the case when $c=1/2$ and Figure 1(b) depicts the case when $c=3/2$. The x-axis corresponds to different $q$ values, and the y-axis is the value of $SR$. The black solid line connects the values of $SR(q\cdot\mathbf{Q}_{0})$, while the solid points represent the proposed statistics $\widehat{SR}(q\cdot\mathbf{Q}_{0})$ in Theorem 2.7.
It is evident that the value of $\widehat{SR}(q\cdot\mathbf{Q}_{0})$ matches the value of $SR(q\cdot\mathbf{Q}_{0})$ very well. As $q$ increases, $SR(q\cdot\mathbf{Q}_{0})$ initially rises and then falls, indicating the presence of an optimal $q$ that maximizes $SR$. With our proposed statistics in Theorem 2.7, we can replicate this trend without requiring knowledge of the true population matrix ${\bm{\Sigma}}$. Furthermore, the close match between $\widehat{SR}(q\cdot\mathbf{Q}_{0})$ and $SR(q\cdot\mathbf{Q}_{0})$ suggests that our estimator can be used to determine the best regularization matrix $\mathbf{Q}$ from a deterministic candidate set, such that $SR(\mathbf{Q})$ achieves the maximum value over the candidate set too.
@@ 4.1.1 Asymptotics with increasing sample size
In this section, we aim to investigate the effect of varying sample size $n$ and data dimension $p$ on our SR predictions, while keeping all other experimental settings consistent with previous settings. We continue to set the parameter $c=p/n$ at either $1/2$ or $3/2$, and focus on observing changes across different $n$ values, specifically $500,1000,1500$, and $2000$. For different values of $n$ and $p$, we still generate and then fix the population covariance matrix ${\bm{\Sigma}}_{0}\in\mathbb{R}^{p\times p}$, the mean vector ${\bm{\mu}}_{0}\in\mathbb{R}^{p}$ and the regularization matrix $\mathbf{Q}_{0}\in\mathbb{R}^{p\times p}$ in the same procedure as before.
Following the same procedure as described above in Section 4.1, for each $(n,p)$ and each given $q$, we obtain $100$ values of $SR(q\cdot\mathbf{Q}_{0})$ and $\widehat{SR}(q\cdot\mathbf{Q}_{0})$. For clear illustration, we let $SR_{b}(q\cdot\mathbf{Q}_{0};n,p)$ and $\widehat{SR}_{b}(q\cdot\mathbf{Q}_{0};n,p)$ denote the values obtained in the $b$-th independent trial ($b\in[100]$) under the pair $(n,p)$. The outcomes are subsequently illustrated in Figure 2.
Figure 2: Simulation results with increasing $n$. Figures 2(a)-2(d) and Figures 2(e)-2(h) respectively correspond to $c=p/n=1/2$ and $c=p/n=3/2$. In Figures 2(a) and 2(e), $y$-axis shows $\sum_{b=1}^{100}(SR_{b}(q\cdot\mathbf{Q}_{0};n,p)-\widehat{SR}_{b}(q\cdot\mathbf{Q}_{0};n,p))^{2}/100$ for different $q$’s. In Figures 2(b) and 2(f), $y$-axis shows $\sum_{b=1}^{100}(SR_{b}(q\cdot\mathbf{Q}_{0};n,p)/\widehat{SR}_{b}(q\cdot\mathbf{Q}_{0};n,p)-1)^{2}/100$ for different $q$’s. In Figures 2(c) and 2(g), $y$-axis gives the boxplot of $\operatorname{argmax}_{q}SR_{b}(q\cdot\mathbf{Q}_{0};n,p)-\operatorname{argmax}_{q}\widehat{SR}_{b}(q\cdot\mathbf{Q}_{0};n,p)$ across the 100 trials for different $n$’s. In Figures 2(d) and 2(h), $y$-axis displays the boxplot of $\max_{q}SR_{b}(q\cdot\mathbf{Q}_{0};n,p)-\max_{q}\widehat{SR}_{b}(q\cdot\mathbf{Q}_{0};n,p)$ across the 100 trials for different $n$’s.
Figure 2 shows that as we increase the sample size $n$, the mean-squared-error (MSE) of $SR(q\cdot\mathbf{Q}_{0};n,p)-\widehat{SR}(q\cdot\mathbf{Q}_{0};n,p)$ (difference) and $SR(q\cdot\mathbf{Q}_{0};n,p)/\widehat{SR}(q\cdot\mathbf{Q}_{0};n,p)-1$ (ratio) starts to get smaller for each $q$ value. So we verified our theory that when $n$ grows to infinity, the estimated value $\widehat{SR}(q\cdot\mathbf{Q}_{0};n,p)$ converges to the value $SR(q\cdot\mathbf{Q}_{0};n,p)$ almost surely, and they are ratio consistent. Additionally, we examine how the optimal $q$ values and the maximum Sharpe ratios of $SR(q\cdot\mathbf{Q}_{0};n,p)$ and $\widehat{SR}(q\cdot\mathbf{Q}_{0};n,p)$ can differ, and Figure 2 clearly demonstrates that as sample size $n$ grows, this discrepancy diminishes, lending further support to our theory.
@@ 4.2 Efficient frontiers
We will use the regularization matrix $\mathbf{Q}=0.2\,\mathbf{Q}_{0}$, where $\mathbf{Q}_{0}$ is defined in Section 4.1. The population covariance matrix ${\bm{\Sigma}}$ and the mean $\mathbf{r}={\bm{\mu}}$ (since the risk-free rate $r_{0}$ is assumed to be zero) we use here are slightly different from those used in Section 4.1. We first generate a vector ${\bm{\xi}}_{3}\in\mathbb{R}^{p}$ whose elements are independently distributed from $\Gamma(1,1)$. Then we let ${\bm{\Sigma}}={\bm{\Sigma}}_{3}={\rm diag}(\lambda_{1},\dots,\lambda_{p})+2\cdot{\bm{1}}{\bm{1}}^{\top}+{\bm{\xi}}_{3}{\bm{\xi}}_{3}^{\top}$. In addition, the mean vector ${\bm{\mu}}$ here has two choices: ${\bm{\mu}}_{3}={\bm{\mu}}_{0}+2\cdot{\bm{1}}$ and ${\bm{\mu}}_{4}={\bm{\mu}}_{0}+2\cdot{\bm{1}}+{\bm{\xi}}_{3}$, where ${\bm{\mu}}_{0}$ is again defined in Section 4.1. The matrix ${\bm{\Sigma}}_{3}$ simulates a covariance matrix with two factors. In the case of ${\bm{\mu}}={\bm{\mu}}_{3}={\bm{\mu}}_{0}+2\cdot{\bm{1}}$, besides betting on individual stock means ${\bm{\mu}}_{0}$, the mean vector also aligns with the market factor ${\bm{1}}$. In the other case of ${\bm{\mu}}={\bm{\mu}}_{4}={\bm{\mu}}_{0}+2\cdot{\bm{1}}+{\bm{\xi}}_{3}$, the mean vector additionally invests upon the second factor ${\bm{\xi}}_{3}$. Note that we do not need to consider the case $\mathbf{r}={\bm{\mu}}={\bm{\mu}}_{0}$ here because if the target return is $\mu_{0}$, the optimal portfolio can be equivalently obtained by setting $\mathbf{r}={\bm{\mu}}_{3}={\bm{\mu}}_{0}+2\cdot{\bm{1}}$ with the updated target return $\mu_{0}+2$, since $\mathbf{w}^{\top}{\bm{\mu}}_{3}=\mathbf{w}^{\top}{\bm{\mu}}_{0}+2(\mathbf{w}^{\top}{\bm{1}})=\mu_{0}+2$.
@@
The following procedure is similar to Section 4.1. For each $(n,p)$, we (1) specify $\mathbf{Q}=0.2\,\mathbf{Q}_{0}$, ${\bm{\Sigma}}={\bm{\Sigma}}_{3}$, and ${\bm{\mu}}={\bm{\mu}}_{3}$ or ${\bm{\mu}}_{4}$; (2) generate 100 independent random matrices $\mathbf{R}\in\mathbb{R}^{n\times p}$ with i.i.d rows following $\mathcal{N}({\bm{\mu}},{\bm{\Sigma}})$; (3) for each given $\mathbf{R}$, compute the sample covariance matrix $\widehat{\bm{\Sigma}}=\frac{(\mathbf{R}-{\bm{1}}_{n}{\bm{\mu}}^{\top})^{\top}(\mathbf{R}-{\bm{1}}_{n}{\bm{\mu}}^{\top})}{n}$; (4) with each $\widehat{\bm{\Sigma}}$, solve the optimization problem in (4.1), for $\mu_{0}$ ranging from $0.2$ to $6$ with the increment of $0.2$; (5) for each $\mu_{0}$, calculate the corresponding values of $\sigma_{0}$ and $\widehat{\sigma}$, whose formula is provided in Theorem 3.2.
According to Theorem 3.2, the volatility of the optimal portfolio $\sigma_{0}$ is defined as $\sigma_{0}^{2}=\mathbf{w}^{*}{\bm{\Sigma}}\mathbf{w}^{*}$, and the corresponding estimator is $\widehat{\sigma}^{2}=\mathbf{w}^{*}\widehat{\bm{\Sigma}}\mathbf{w}^{*}/(1-c/p\cdot{\mathrm{tr}}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1})^{2}$. We conduct simulations to confirm that $\widehat{\sigma}_{0}$ matches $\sigma_{0}$ well for each given $\mu_{0}$. The results displayed in Figure 3 are the average over the 100 values of $\sigma_{0}$ and $\widehat{\sigma}$ for different $\mu_{0}$.
@@
Figure 3: Efficient frontiers of $\mathbf{w}^{*}$ from (4.1). The x-axis is the volatility level, and the y-axis is the target return $\mu_{0}$. The solid line characterizes the curve of $(\mu_{0},\sigma_{0})$, while the solid points represent the points in the curve of $(\mu_{0},\widehat{\sigma})$, where $\widehat{\sigma}$ is given in Theorem 3.2.
@@
As shown in Figure 3, we observe that $\widehat{\sigma}$ estimates $\sigma_{0}$ very accurately for various $\mu_{0}$ and different $c$. As $\mu_{0}$ increases, we notice that the volatility $\sigma_{0}$ initially decreases and then increases, actually forming a hyperbola. Remarkably, our estimation $\widehat{\sigma}$ also closely depicts this curve, even without access to the population covariance matrix.

### v3

@@ 4 Numerical Experiments
In this section, we perform numerical experiments to validate our theoretical findings. In Section 4.1, we provide the basic settings for SR estimation and verify Theorem 2.7, especially the asymptotic behavior as $n$ increases. In Section 4.2, we verify our estimation of the efficient frontier in Theorem 3.2. Some additional simulations are given in Appendix A. Appendix A.1 provides comparisons on different choices of the population covariance ${\bm{\Sigma}}$, the mean vector of returns ${\bm{\mu}}$ and the regularization $\mathbf{Q}$, while Appendix A.2 gives simulations on unknown ${\bm{\mu}}$. In all the simulations, we set $r_{0}=0$ for simplicity so that $\mathbf{r}={\bm{\mu}}$.
@@ 4.1 Sharpe ratio estimation
We present the simulation results under the basic settings. We first fix the sample size as $n=1500$ and consider two dimensions: $p=750$ ($c=1/2$) and $p=2250$ ($c=3/2$). These cases correspond to values of $c$ less than and greater than 1. To simulate financial returns, we construct the population covariance ${\bm{\Sigma}}_{0}$, mean vector ${\bm{\mu}}_{0}$, and the matrix $\mathbf{Q}_{0}$ as follows:
@@
    The population covariance ${\bm{\Sigma}}={\bm{\Sigma}}_{0}={\rm diag}(\lambda_{1},\dots,\lambda_{p})+2{\bm{1}}{\bm{1}}^{\top}$, where $\{\lambda_{i}\}_{i=1}^{p}$ are generated from a truncated $\Gamma^{-1}(1,1)$ distribution, truncated with the interval $[0.01,9]$, and then ranked in decreasing order. This particular construction of ${\bm{\Sigma}}$ follows from the factor model. Specifically, the covariance consists of the market systematic risk, where the vector ${\bm{1}}$ represents a market factor introducing a spike in covariance spectrum, and the specific risk, where $\lambda_{i}$ represents the individual residual risk of each asset.
@@
    The regularization matrix is $\mathbf{Q}=q\cdot\mathbf{Q}_{0}$, where $\mathbf{Q}_{0}={\rm diag}(3,...,3,1,...,1)$ with $p/2$ numbers of $3$ and $1$ entries. By structuring $\mathbf{Q}_{0}$ in this manner, we categorize stocks into high and low volatility regimes, capturing some weak knowledge of the covariance structure. We will vary the values of $q$ in order to verify Theorem 2.7.
In the simulation, for each pair of $(n,p)$ considered, the parameters ${\bm{\Sigma}}_{0}\in\mathbb{R}^{p\times p}$, ${\bm{\mu}}_{0}\in\mathbb{R}^{p}$, and $\mathbf{Q}_{0}\in\mathbb{R}^{p\times p}$ are generated, and they will remain fixed throughout the simulation procedure. Then we generate iid Gaussian random vector $\mathbf{R}_{i}$ ($i\in[n]$) with mean ${\bm{\mu}}={\bm{\mu}}_{0}$ and variance ${\bm{\Sigma}}={\bm{\Sigma}}_{0}$. The observed random matrix is $\mathbf{R}=(\mathbf{R}_{1},\mathbf{R}_{2},...,\mathbf{R}_{n})^{\top}\in\mathbb{R}^{n\times p}$. In this simulation, we assume the mean vector ${\bm{\mu}}$ is known, and thus $\mathbf{X}=\mathbf{R}-{\bm{1}}_{n}{\bm{\mu}}_{0}^{\top}$ is also well observed. Recall that $\widehat{\bm{\Sigma}}=\mathbf{X}^{\top}\mathbf{X}/n$, and $SR(\mathbf{Q})=\frac{{\bm{\mu}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\mu}}}{\sqrt{{\bm{\mu}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\mu}}}}$. We compare the true $SR(\mathbf{Q})$ and the estimator $\widehat{SR}(\mathbf{Q})$ in Theorem 2.7 which is
@@
The Sharpe computation is then carried out as follows for each pair of $(n,p)$. We repeat the generation of Gaussian random matrices $\mathbf{R}\in\mathbb{R}^{n\times p}$ for $1000$ independent times. For each repetition, we vary the value of $q$ to compute and compare $SR(q\cdot\mathbf{Q}_{0})$ and $\widehat{SR}(q\cdot\mathbf{Q}_{0})$. If $c<1$, we vary $q$ from (1:30)/5(1:30)/5( 1 : 30 ) / 5, while if $c>1$, we vary $q$ from (1:30)/1.5(1:30)/1.5( 1 : 30 ) / 1.5. For each given $q$, we will obtain 1000 different values of $SR(q\cdot\mathbf{Q}_{0})$ and $\widehat{SR}(q\cdot\mathbf{Q}_{0})$.
In Figure 1, the $x$-axis represents the different values of $q$, while the $y$-axis gives the Sharpe ratio. The simulation results are presented as the average of $SR(q\cdot\mathbf{Q}_{0})$ and $\widehat{SR}(q\cdot\mathbf{Q}_{0})$ over $1000$ independent trials.
Figure 1: Simulation results in the basic settings. Figure 1(a) shows the case when $c=1/2$ and Figure 1(b) depicts the case when $c=3/2$. The x-axis corresponds to different $q$ values, and the y-axis is the value of $SR$. The black solid line connects the values of $SR(q\cdot\mathbf{Q}_{0})$, while the points represent the proposed statistics $\widehat{SR}(q\cdot\mathbf{Q}_{0})$ in Theorem 2.7.
It is evident that the value of $\widehat{SR}(q\cdot\mathbf{Q}_{0})$ matches the value of $SR(q\cdot\mathbf{Q}_{0})$ very well. As $q$ increases, $SR(q\cdot\mathbf{Q}_{0})$ initially rises and then falls, indicating the presence of an optimal $q$ that maximizes $SR$. With our proposed statistics in Theorem 2.7, we can replicate this trend without requiring knowledge of the true population matrix ${\bm{\Sigma}}$. Furthermore, the close match between $\widehat{SR}(q\cdot\mathbf{Q}_{0})$ and $SR(q\cdot\mathbf{Q}_{0})$ suggests that our estimator can be used to determine the best regularization matrix $\mathbf{Q}$ from a deterministic candidate set, such that $SR(\mathbf{Q})$ achieves the maximum over the candidate set too.
@@ 4.1.1 Asymptotics with increasing sample size
In this section, we aim to investigate the effect of varying $n$ and $p$ on our SR predictions, while keeping all other experimental settings consistent with previous settings. We continue to set the parameter $c=p/n$ at either $1/2$ or $3/2$, and focus on observing changes across different $n$ values, specifically $500,1000,1500$, and $2000$. Following the same procedure as described above in Section 4.1, for each $(n,p)$ and each given $q$, we obtain $1000$ values of $SR(q\cdot\mathbf{Q}_{0})$ and $\widehat{SR}(q\cdot\mathbf{Q}_{0})$. For clear illustration, we let $SR_{b}(q\cdot\mathbf{Q}_{0};n,p)$ and $\widehat{SR}_{b}(q\cdot\mathbf{Q}_{0};n,p)$ denote the values obtained in the $b$-th independent trial ($b\in[1000]$) under the pair $(n,p)$. The outcomes are subsequently illustrated in Figure 2.
Figure 2: Simulation results with increasing $n$. Figures 2(a)-2(d) and Figures 2(e)-2(h) respectively correspond to $c=p/n=1/2$ and $c=p/n=3/2$. In Figures 2(a) and 2(e), $y$-axis shows $\sum_{b=1}^{1000}(SR_{b}(q\cdot\mathbf{Q}_{0};n,p)-\widehat{SR}_{b}(q\cdot\mathbf{Q}_{0};n,p))^{2}/1000$ for different $q$’s. In Figures 2(b) and 2(f), $y$-axis shows $\sum_{b=1}^{1000}(SR_{b}(q\cdot\mathbf{Q}_{0};n,p)/\widehat{SR}_{b}(q\cdot\mathbf{Q}_{0};n,p)-1)^{2}/1000$ for different $q$’s. In Figures 2(c) and 2(g), $y$-axis gives the boxplot of $\operatorname{argmax}_{q}SR_{b}(q\cdot\mathbf{Q}_{0};n,p)-\operatorname{argmax}_{q}\widehat{SR}_{b}(q\cdot\mathbf{Q}_{0};n,p)$ across the 1000 trials for different $n$’s. In Figures 2(d) and 2(h), $y$-axis displays the boxplot of $\max_{q}SR_{b}(q\cdot\mathbf{Q}_{0};n,p)-\max_{q}\widehat{SR}_{b}(q\cdot\mathbf{Q}_{0};n,p)$ across the 1000 trials for different $n$’s.
Figure 2 shows that as we increase the sample size $n$, the mean-squared-errors (MSE) of $SR(q\cdot\mathbf{Q}_{0};n,p)-\widehat{SR}(q\cdot\mathbf{Q}_{0};n,p)$ (difference) and $SR(q\cdot\mathbf{Q}_{0};n,p)/\widehat{SR}(q\cdot\mathbf{Q}_{0};n,p)-1$ (ratio) get smaller for each $q$ value. So we verified our theory that when $n\to\infty$, the estimated value $\widehat{SR}(q\cdot\mathbf{Q}_{0};n,p)$ converges to the value $SR(q\cdot\mathbf{Q}_{0};n,p)$ almost surely, and they are ratio consistent. Additionally, we examine how the optimal $q$ values and the maximal Sharpe ratios of $SR(q\cdot\mathbf{Q}_{0};n,p)$ and $\widehat{SR}(q\cdot\mathbf{Q}_{0};n,p)$ can differ. Figure 2 clearly demonstrates that as $n$ grows, this discrepancy diminishes, lending further support to our theory.
@@ 4.2 Efficient frontiers
We will use the regularization matrix $\mathbf{Q}=0.2\mathbf{Q}_{0}$, where $\mathbf{Q}_{0}$ is defined in Section 4.1. The population covariance matrix ${\bm{\Sigma}}$ and the mean $\mathbf{r}={\bm{\mu}}$ (since the risk-free rate $r_{0}$ is assumed to be zero) we use here are slightly different from those used in Section 4.1. We first generate a vector ${\bm{\xi}}_{3}\in\mathbb{R}^{p}$  [^4] Notations ${\bm{\xi}}_{1},{\bm{\xi}}_{2},{\bm{\mu}}_{1},{\bm{\mu}}_{2},{\bm{\Sigma}}_{1},{\bm{\Sigma}}_{2},\mathbf{Q}_{1},\mathbf{Q}_{2}$ are used in Appendix A.1. whose elements are independently distributed from $\Gamma(1,1)$. Then we let ${\bm{\Sigma}}={\bm{\Sigma}}_{3}={\rm diag}(\lambda_{1},\dots,\lambda_{p})+2\cdot{\bm{1}}{\bm{1}}^{\top}+{\bm{\xi}}_{3}{\bm{\xi}}_{3}^{\top}$. The matrix ${\bm{\Sigma}}_{3}$ simulates a covariance matrix with two factors. In addition, the mean vector ${\bm{\mu}}$ here has two choices: ${\bm{\mu}}_{3}=p^{\frac{1}{4}}{\bm{\mu}}_{0}+2\cdot{\bm{1}}$ and ${\bm{\mu}}_{4}={\bm{\mu}}_{0}+2\cdot{\bm{1}}+{\bm{\xi}}_{3}$, where ${\bm{\mu}}_{0}$ is again defined in Section 4.1. We design the vectors ${\bm{\mu}}_{3}$ and ${\bm{\mu}}_{4}$ such that $\mathcal{A}_{rr}$ becomes unbounded when ${\bm{\mu}}={\bm{\mu}}_{3}$, while it remains bounded when ${\bm{\mu}}={\bm{\mu}}_{4}$.
@@
The following procedure is similar to Section 4.1. For each $(n,p)$, we (1) specify $\mathbf{Q}=0.2\mathbf{Q}_{0}$, ${\bm{\Sigma}}={\bm{\Sigma}}_{3}$, and ${\bm{\mu}}={\bm{\mu}}_{3}$ or ${\bm{\mu}}_{4}$; (2) generate 1000 independent random matrices $\mathbf{R}\in\mathbb{R}^{n\times p}$ with i.i.d rows following $\mathcal{N}({\bm{\mu}},{\bm{\Sigma}})$; (3) for each given $\mathbf{R}$, compute the sample covariance matrix $\widehat{\bm{\Sigma}}=\frac{(\mathbf{R}-{\bm{1}}_{n}{\bm{\mu}}^{\top})^{\top}(\mathbf{R}-{\bm{1}}_{n}{\bm{\mu}}^{\top})}{n}$; (4) with each $\widehat{\bm{\Sigma}}$, solve the optimization problem in (4.1), for $\mu_{0}$ ranging from $0.2$ to $6$ with the increment of $0.2$; (5) for each $\mu_{0}$, calculate the corresponding values of $\sigma_{0}$ and $\widehat{\sigma}$, whose formula is provided in Theorem 3.2. According to Theorem 3.2, the volatility of the optimal portfolio $\sigma_{0}$ is defined as $\sigma_{0}^{2}=\mathbf{w}^{*}{\bm{\Sigma}}\mathbf{w}^{*}$, and the corresponding estimator is $\widehat{\sigma}^{2}=\mathbf{w}^{*}\widehat{\bm{\Sigma}}\mathbf{w}^{*}/(1-c/p\cdot{\mathrm{tr}}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1})^{2}$. We conduct simulations to confirm that $\widehat{\sigma}_{0}$ matches $\sigma_{0}$ well for each given $\mu_{0}$. The results displayed in Figure 3 are the average over the 1000 values of $\sigma_{0}$ and $\widehat{\sigma}$ for different $\mu_{0}$’s.
@@
Figure 3: Efficient frontiers of $\mathbf{w}^{*}$ from (4.1). The x-axis is the volatility level, and the y-axis is the target return $\mu_{0}$. The solid line characterizes the curve of $(\mu_{0},\sigma_{0})$, while the points represent $(\mu_{0},\widehat{\sigma})$, where $\widehat{\sigma}$ is given in Theorem 3.2.
@@
As shown in Figure 3, we observe that $\widehat{\sigma}$ estimates $\sigma_{0}$ very accurately for various $\mu_{0}$ and different $c$. As $\mu_{0}$ increases, we notice that the volatility $\sigma_{0}$ initially decreases and then increases, actually forming a hyperbola. Remarkably, $\widehat{\sigma}$ closely depicts this curve.